OpenStudy (phebe):

MEDALSSS!!! 13. A ball is thrown into the air. The height in feet of the ball can be modeled by the equation h = -16t2 + 20t + 6 where t is the time in seconds, the ball is in the air. When will the ball hit the ground? How high will the ball go?

OpenStudy (gaberen):

Set h(t) = 0 to find when the ball will hit the ground You should get t = 1.5 (t has to be positive) The max height occurs at the vertex at the parabola which is on the axis of symmetry. Axis of symmetry = \[-b/2a=-20/-32=5/8\]. Plug this value into h(t) to obtain the vertex which is (0.625,12.25).

OpenStudy (phebe):

okk

OpenStudy (anonymous):

You can see that you evidently have a quadratic equation so may use the quadratic formula to solve for the "positive' value

OpenStudy (phebe):

so i divide or multiply

OpenStudy (phebe):

@Johnbc

OpenStudy (phebe):

@gaberen

OpenStudy (anonymous):

Divide or multiply what?

OpenStudy (dan815):

use the quanratic formula, or complete the square , set h=0 and u have to solve for t

OpenStudy (dan815):

because your motion is described by a negative parabola, so there will be 2 points when height =0 once when its launched and when it hits the ground again

OpenStudy (dan815):

you want that 2nd longer time

OpenStudy (dan815):